determine the truth value of each of the following truth values

if 2+3=5 then 2x3=6



1
Expert's answer
2021-03-18T09:06:46-0400

Let determine the truth value of the statement "if "2+3=5" then "2\times3=6"". Since the statements ""2+3=5"" and ""2\times3=6"" are true, their implication, that is the statement "if "2+3=5" then "2\times3=6"" is also true.